OpenStudy (anonymous):

how do i find tan 67.5 with the tangant half angle formula???

OpenStudy (asnaseer):

what is 2 times 67.5?

OpenStudy (anonymous):

135

OpenStudy (anonymous):

Half-Angle formulas sin s/2=+-sqrt((1-cos s)/2), (note:+ or - answer depends on value of sin s/2 not sin s.) cos s/2=+-sqrt((1+cos s)/2), (note:+ or - answer depends on value of cos s/2 not cos s.) tan s/2=sin s/(1+cos s) .. sin 67.5=sin 135/2=sqrt((1-cos 135)/2) =sqrt((1+√2/2)/2)=sqrt((2+√2)/4) =sqrt(2+√2)/2 .. cos 67.5=cos 135/2=sqrt((1+cos 135)/2 =sqrt((1+(-√2/2))/2)=sqrt((1-√2/2))/2) =sqrt((2-√2)/4) =sqrt(2-√2)/2 .. tan 67.5=tan 135/2=sin 135/(1+cos 135) =(√2/2)/1+(-√2/2) =√2/2)/1-(√2/2)) =√2/(2-√2) Note: ( + answers selected for sin & cos because 67.5º is in quadrant I where sin & cos are positive) http://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.444897.html

OpenStudy (asnaseer):

@Lulu212 now use the fact that:\[\tan(\theta)=-\tan(180-\theta)\]to see what you get for \(\tan(135)\)

OpenStudy (anonymous):

@asnaseer he asks for the tangant half angle formula???

OpenStudy (asnaseer):

I'm getting there @Yahoo! - be patient. :)

OpenStudy (anonymous):

OK!..

OpenStudy (asnaseer):

@Lulu212 can you answer my question above ^^^ ?

OpenStudy (anonymous):

180 - theta = 135 find theta?

OpenStudy (asnaseer):

not quiet:\[\tan(135)=-\tan(180-135)=-\tan(45)=?\]@Lulu212 do you know what the \(\tan(45)\) is equal to?

OpenStudy (asnaseer):

@Lulu212 I can only help you here if you respond to my questions.